TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint8x2007x06x04x08x53

Cooldown Sector 78

The commissioning programme on Q4, Q5, D2 and the associated correctors continued during this week. While the Q4, Q5 and D2 commissioning was completed, the commissioning of the 120 A correctors associated to these magnets could not be completed: some additional tuning of the software in the FGC remains to be done.

Yesterday (Thursday May30th) Q4, Q5, D2 were taken together to nominal current. At about 80% of the ramp, D2 quenched (5268 A). As expected, this caused the shutdown of the converters in the matching section and during the ramp down Q4 and Q5 also quenched. The data of this event is still being analysed.

The Electrical Quality Assurance team qualified the focusing and the defocusing quadrupole circuits and the matching section quadrupole Q6 at Point 7.

All the DFBs of the sector have been equipped with dry-air bags to ensure no condensation forms on the current leads and their chimneys.

Unfortunately, on Friday (June 1st) morning, the safety chain of cabinet was accidentally broken and the sector was disconnected from the pumping group. The cryo team is now busy recovering the situation to re-establish nominal temperature on the magnets for Monday morning.

Comments:

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-06-04 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login